# LSST DESC Collaboration

### v1

@@ 1 Introduction
Milky Way satellite dwarf galaxies and halo star clusters (which we collectively refer to as Milky Way “satellites”) are detected as arcminute-scale statistical over-densities of individually resolved stars (e.g., Willman et al. 2011). More than 65 confirmed and candidate dark-matter-dominated satellite galaxies have been detected around the Milky Way to date (e.g., Pace 2024, and references therein). When correcting for observational completeness and accounting for theoretical uncertainties, the known population of satellite galaxies is found to agree with predictions from the CDM model (e.g., Kim et al. 2018; Nadler et al. 2020; Manwadkar & Kravtsov 2022). However, observational and theoretical arguments suggest that the current census of Milky Way satellites is incomplete, and recent models predict that the total Milky Way satellite galaxy populations consists of $\sim$100–300 systems (e.g., Tollerud et al. 2008; Koposov et al. 2008; Walsh et al. 2008; Jethwa et al. 2018; Newton et al. 2018; Kim et al. 2018; Drlica-Wagner et al. 2020; Nadler et al. 2020; Manwadkar & Kravtsov 2022; Nadler et al. 2024). In addition, more than a dozen ultra-faint, compact stellar systems have been discovered in the outer Milky Way halo with a physical nature that is unclear (e.g., Koposov et al. 2008; Balbinot et al. 2013; Kim et al. 2015; Mau et al. 2019; Cerny et al. 2021, 2023a, 2023b; Smith et al. 2024). These systems may be faint star clusters that formed in dwarf galaxies that were accreted and disrupted (e.g., Malhan et al. 2022), or they may be an extension of the dwarf galaxy population into the small, hyper-faint regime (e.g., Manwadkar & Kravtsov 2022; Errani et al. 2023).
@@
Thus far, the detection of ultra-faint Milky Way satellites has been limited by the sensitivity and sky coverage of observational surveys (e.g., Koposov et al. 2008; Walsh et al. 2008; Drlica-Wagner et al. 2020; Homma et al. 2023). The upcoming Legacy Survey of Space and Time (LSST) that will be performed by the NSF-DOE Vera C. Rubin Observatory is expected to greatly advance the frontier of ultra-faint satellite discovery. LSST is an optical/near-infrared survey that is expected to reach a coadded imaging depth of $r\sim 27.5$ mag (${\rm S/N}=5$, point-like sources) by the end of its 10-year mission (Ivezić et al. 2019). This sensitivity is expected to enable the detection of resolved stellar systems around the Milky Way with surface brightnesses as faint as $\mu_{V}\sim 32$ mag arcsec-2 over a sky area covering $\sim$20,000 deg2 (Tollerud et al. 2008; Hargis et al. 2014). Furthermore, the location of the LSST footprint in the southern hemisphere provides additional sensitivity to satellites of the Magellanic Clouds (e.g., Kallivayalil et al. 2018; Patel et al. 2020). LSST is expected to take a significant step toward completing the census of the faintest known galaxies ($M_{V}\sim 0$ mag) out to the virial radius ($\sim$300 kpc) of the Milky Way (e.g., Newton et al. 2018; Nadler et al. 2019).
@@ 2.1 Simulated LSST Data
The starting point for our analysis is simulated data from LSST DESC DC2, an end-to-end simulated sky survey developed in preparation for extragalactic analyses of LSST data (Abolfathi et al. 2021; LSST DESC et al. 2022). Input object samples were assembled from an N-body simulation of large-scale structure populated with galaxies (i.e., the CosmoDC2 synthetic galaxy catalog based on the Outer Rim simulations; Korytov et al. 2019; Heitmann et al. 2019) and a stellar population simulated to mimic the Milky Way bulge, disk, and halo (i.e., based on galfast; Jurić et al. 2008). DC2 consists of five years of simulated LSST observations generated with then-current estimates of the survey cadence and environmental conditions (minion_1016; Jones et al. 2015). Realistic images were simulated by the imSim software, [^3] https://github.com/LSSTDESC/imSim/releases/tag/v0.6.2 and processed with the LSST Science Pipelines to correct for simulated instrument signatures and produce object catalogs. The simulated DC2 sky survey includes six optical bands, ugrizy, covering $\sim$300 deg2 of the LSST wide-fast-deep (WFD) footprint, as well as a deep drilling field (DDF) of approximately 1 deg2. The median ${\rm S/N}=5$ magnitude limits for point-like sources in the DC2 WFD simulations are $g=27.0$ mag and $r=26.8$ (Figure 1).
The DC2 simulations have no throughput variation over the focal plane or between visits, so joint photometric and astrometric calibration across visits is not performed. Consequently, the standard passbands for the LSST filters are simply the total throughputs used in the simulations and were derived by the Rubin systems engineering team. [^4] https://github.com/lsst/throughputs/tree/DC2production We used these same throughputs to generate the stellar populations of simulated dwarf galaxies as described in Section 2.4.
Figure 1: The simulated LSST DESC DC2 footprint covers $\sim$300 deg2 of the high Galactic latitude sky. Maps of the ${\rm S/N}=5$ magnitude limit for point-like sources in the $g$-band (left) and $r$-band (right) show the uniformity and depth of the WFD portion of DC2. The median ${\rm S/N}=5$ magnitude limits for point-like sources are $g=27.0$ mag and $r=26.8$ mag.
@@ 2.2 Data Access
We accessed and analyzed the DC2 data through the RSP hosted on the Interim Data Facility (IDF) in the context of the first data preview (DP0.1) released in the period leading up to the start of LSST (Guy et al. 2023). The goals of DP0.1 were to serve as an early integration test of the LSST Science Pipelines with the RSP, and to enable members of the scientific community to begin early preparations for science with LSST. DP0.1 was hosted on the IDF and access was provided through the RSP, which includes a set of integrated web-based applications, services, and tools to query, visualize, subset, and analyze LSST data (O’Mullane et al. 2021). The $\sim$300 deg2 WFD data set from DC2 was ingested and adopted as the primary data set for DP0.1.555https://dp0-1.lsst.io/data-products-dp0-1 In addition to the simulated output catalogs, DP0.1 contains a truth matched catalog to access properties of the astronomical objects used as input to DC2. All analyses presented in this paper were performed using the RSP on the IDF, and a discussion of how we optimized our analysis to run on the RSP is provided in Appendix A.
@@ 2.3 Catalog Characterization
To quantify star/galaxy separation efficiency, we selected a $\sim$3 deg2 region of the DC2 WFD footprint centered on $({\rm R.A.,Dec.})=(62\,\deg,-37\,\deg)$. Within this region, we selected a set of well-measured objects with truth matches and $g$- and $r$-band magnitude error, $\sigma_{g,r}<0.2$ (S/N $\gtrsim 5$). We performed a simple star/galaxy separation based on the measured EXTENDEDNESS parameter (Bosch et al. 2018; LSST DESC et al. 2022). The EXTENDEDNESS is a boolean classification variable defined from the difference between the point spread function (PSF) magnitude and the CModel magnitude for objects. Point-like sources are have $\texttt{EXTENDEDNESS}=0$, while extended galaxies have $\texttt{EXTENDEDNESS}=1$ (LSST DESC et al. 2022). To characterize the detection efficiency, we also tracked all stars in the region that were truth matched regardless of whether they passed our $\sigma_{g,r}<0.2$ threshold. We binned these populations in true $r$-band magnitude to derive the stellar detection efficiency, stellar classification efficiency, and the simultaneous combination of the two (Figure 2). We find a steep drop in stellar classification efficiency by $r\sim 26$ mag.
@@
To derive a model for the photometric uncertainty as a function of object magnitude and survey depth, we used the measured magnitude uncertainties of stars in combination with ${\rm S/N}=5$ magnitude limit depth maps generated from the input images (Figure 1). The magnitude limit maps were generated from the simulated observing properties of the DC2 simulations using the suprême package [^6] https://github.com/LSSTDESC/supreme/. We compute the median of the logarithm of the $r$-band magnitude uncertainty as a function of the difference between the true input $r$-band magnitude of the simulated object and the value of the magnitude limit map at the location of the object, $\Delta r=r-{\rm maglim}(r)$ (Figure 2). This model for the expected magnitude uncertainty as a function of magnitude and survey magnitude limit is applied to the simulated satellite member stars to generate realistic photometric scatter in magnitude and color (Section 2.4).
Figure 3: Simulated LSST color–magnitude diagram of stars associated with a simulated Milky Way satellite (red points; $M_{V}=-0.3$ mag, $D=91$ kpc) and the density of simulated objects from DESC DC2 that are classified as stars (gray background). The left panel uses a measured star/galaxy classification based on the EXTENDEDNESS parameter while the right panel assumes a perfect star/galaxy separation. The red solid line shows our isochrone selection region, and the dashed red line indicates the magnitude cut at $g<26$ mag and $r<26$ mag that is applied in our analysis. The population of faint ($g\gtrsim 25$ mag), blue ($g-r\lesssim 0.25$) background galaxies that are mis-classified as stars presents a significant contamination in searches for faint Milky Way satellites.
@@ 2.4 Simulated Satellites
The simulated LSST bandpasses used for DC2 were integrated into the Padova CMD interface to generate photometry consistent with DC2.777http://stev.oapd.inaf.it/cgi-bin/cmd The initial masses of satellite member stars were drawn from a Chabrier (2001) initial mass function (IMF), which has been found to be a reasonable description of known satellite galaxies (Simon 2019). These initial masses were used to assign current absolute magnitudes from a Marigo et al. (2017) isochrone. When sampling from the IMF, the lower mass bound was set to the hydrogen-burning limit of $0.08M_{\odot}$ and the upper bound was set by the star with the largest initial mass in the evolved isochrone (post-asymptotic giant branch and white dwarf stars are excluded). Using the Marigo et al. (2017) isochrones, we transformed from initial stellar mass to current absolute magnitude in the $g$ and $r$ bands of the DC2 photometric system, and then to apparent magnitudes using the distance modulus of the simulated satellite. Detectability, stellar classification probability, photometric scatter, and photometric measurement uncertainties were assigned to each simulated star based on its true apparent magnitude and the DC2 survey limiting magnitude at the location that the star was injected following the procedures described in Section 2.3. In this way, our simulated stellar catalog mimics the DC2 catalog in detection efficiency, object (mis-)classification, and photometric uncertainty, and thus it can be combined with the DC2 catalog directly.
@@ 3 Analysis
To explore the impact of star/galaxy classification efficiency, we ran the analysis in two different configurations. In the first, the star/galaxy classification was performed using the measured EXTENDEDNESS parameter: objects with $\texttt{EXTENDEDNESS}=0$ were classified as stars, and those with $\texttt{EXTENDEDNESS}=1$ were classified as background galaxies. In the second configuration, the star/galaxy classification instead used the true object class from the simulation inputs. The difference in star/galaxy classification efficiency was also applied when generating simulated satellites—i.e., we generated two separate sets of satellite simulations applying the measured and perfect star/galaxy classification efficiency. Example color–magnitude diagrams of the objects passing these selections for one simulated satellite are shown in Figure 3. Downstream analysis procedures were the same for these two different star/galaxy separation procedures.
@@ 6 Conclusions
There is significant room to improve star/galaxy classification, both with LSST data alone and in combination with upcoming space missions, such as the Nancy Grace Roman Space Telescope (Spergel et al. 2015). Proposed synergies between Rubin and Roman suggest that the high spatial resolution of Roman ($\sim 0.1"$ full-width at half maximum) could efficiently distinguish point sources from extended sources down to 25 AB mag at $5\sigma$ depth using the high-throughput F146 filter (e.g., Han et al. 2023). There also exist novel ideas for improving star/galaxy separation efficiency using multi-wavelength observations and advanced image processing techniques (e.g., Kovács & Szapudi 2015; Sevilla-Noarbe et al. 2018; Muyskens et al. 2022). Satellite detection efficiency can also be improved by utilizing likelihood-based search algorithms (e.g., Bechtol et al. 2015; Drlica-Wagner et al. 2020). These algorithms construct a likelihood function from the product of Poisson probabilities to detect individual stars based upon their spatial positions, measured fluxes, photometric uncertainties, and the local imaging depth, given a model that includes a putative dwarf galaxy and empirical estimation of the local stellar field population. The combination of the simple matched-filter algorithm applied here with a likelihood-based search approach has been found to greatly reduce false positives (Drlica-Wagner et al. 2020).
@@
Software: Astropy (Astropy Collaboration et al. 2013, 2018), HealPy (Zonca et al. 2019) healsparse (Rykoff & Sánchez 2019), Matplotlib (Hunter 2007), Numpy (Harris et al. 2020), Pandas (pandas development team 2020), Phalanx [^14] https://phalanx.lsst.io/, SciPy (Virtanen et al. 2020), simple (Bechtol et al. 2015), ugali (Bechtol et al. 2015; Drlica-Wagner et al. 2020)

### v2

# Predictions for the Detectability of Milky Way Satellite Galaxies and Outer-Halo Star Clusters with the Vera C. Rubin Observatory

@@ 1 Introduction
Milky Way satellite dwarf galaxies and halo star clusters (which we collectively refer to as Milky Way “satellites”) are detected as arcminute-scale statistical over-densities of individually resolved stars (e.g., Willman et al. 2011). More than 65 confirmed and candidate dark-matter-dominated satellite galaxies have been detected around the Milky Way to date (e.g., Pace 2024, and references therein). When correcting for observational completeness and accounting for theoretical uncertainties, the known population of satellite galaxies is found to agree with predictions from the CDM model (e.g., Kim et al. 2018; Nadler et al. 2020; Manwadkar & Kravtsov 2022). However, observational and theoretical arguments suggest that the current census of Milky Way satellites is incomplete, and recent models predict that the total Milky Way satellite galaxy populations consists of $\sim$100–300 systems (e.g., Tollerud et al. 2008; Koposov et al. 2008; Walsh et al. 2008; Jethwa et al. 2018; Newton et al. 2018; Kim et al. 2018; Drlica-Wagner et al. 2020; Nadler et al. 2020; Manwadkar & Kravtsov 2022; Nadler et al. 2024). In addition, more than a dozen ultra-faint, compact stellar systems have been discovered in the outer Milky Way halo with a physical nature that is unclear (e.g., Koposov et al. 2008; Balbinot et al. 2013; Kim et al. 2015; Mau et al. 2019; Cerny et al. 2021, 2023a, 2023b; Smith et al. 2024). These systems may be faint star clusters that formed in dwarf galaxies that were accreted and disrupted (e.g., Malhan et al. 2022), or they may be an extension of the dwarf galaxy population into the small, hyper-faint regime (e.g., Manwadkar & Kravtsov 2022; Errani et al. 2024).
@@
Thus far, the detection of ultra-faint Milky Way satellites has been limited by the sensitivity and sky coverage of observational surveys (e.g., Koposov et al. 2008; Walsh et al. 2008; Drlica-Wagner et al. 2020; Homma et al. 2024). The upcoming Legacy Survey of Space and Time (LSST) that will be performed by the NSF-DOE Vera C. Rubin Observatory is expected to greatly advance the frontier of ultra-faint satellite discovery. LSST is an optical/near-infrared survey that is expected to reach a coadded imaging depth of $r\sim 27.5$ mag (${\rm S/N}=5$, point-like sources) by the end of its 10-year mission (Ivezić et al. 2019). This sensitivity is expected to enable the detection of resolved stellar systems around the Milky Way with surface brightnesses as faint as $\mu_{V}\sim 32$ mag arcsec-2 over a sky area covering $\sim$20,000 deg2 (Tollerud et al. 2008; Hargis et al. 2014). Furthermore, the location of the LSST footprint in the southern hemisphere provides additional sensitivity to satellites of the Magellanic Clouds (e.g., Kallivayalil et al. 2018; Patel et al. 2020). LSST is expected to take a significant step toward completing the census of the faintest known galaxies ($M_{V}\sim 0$ mag) out to the virial radius ($\sim$300 kpc) of the Milky Way (e.g., Newton et al. 2018; Nadler et al. 2019).
@@ 2.1 Simulated LSST Data
The starting point for our analysis is simulated data from LSST DESC DC2, an end-to-end simulated sky survey developed in preparation for extragalactic analyses of LSST data (Abolfathi et al. 2021; LSST DESC et al. 2022). The DC2 simulations aim to capture the characteristics of high-Galactic-latitude LSST observations in a small ( ${\sim}\,300$ deg2) region centered on ${\rm RA},{\rm Dec.}=61.863^{\circ},-35.790^{\circ}$ ($\ell,b=237.34^{\circ},-47.75^{\circ}$). Input object samples were assembled from an N-body simulation of large-scale structure populated with galaxies (i.e., the CosmoDC2 synthetic galaxy catalog based on the Outer Rim simulations; Korytov et al. 2019; Heitmann et al. 2019) and a stellar population simulated to mimic the Milky Way bulge, disk, and halo (i.e., based on galfast; Jurić et al. 2008). Resolved stellar populations from Milky Way halo substructures (i.e., the Magellanic Clouds, other satellite galaxies, and stellar streams) were not included in the DC2 stellar catalog. The fluxes of simulated sources were reddened due to interstellar dust using the extinction maps from Schlegel et al. (1998); however, the effects of reddening are small due to the minimal extinction in the DC2 field, $E(B-V)\lesssim 0.07$. DC2 consists of five years of simulated LSST observations generated with then-current estimates of the survey cadence and environmental conditions (minion_1016; Jones et al. 2015). Realistic images were simulated by the imSim software, [^3] https://github.com/LSSTDESC/imSim/releases/tag/v0.6.2 and processed with the LSST Science Pipelines to correct for simulated instrument signatures and produce object catalogs. The simulated DC2 sky survey includes six optical bands, ugrizy, covering $\sim$ 300 deg2 of the LSST wide-fast-deep (WFD) footprint, as well as a deep drilling field (DDF) of approximately 1 deg2. The median ${\rm S/N}=5$ magnitude limits for point-like sources in the DC2 WFD simulations are $g=27.0$ mag and $r=26.8$ mag (Figure 1). The magnitude limit is estimated from the properties of the simulated DC2 observations (e.g., focal plane geometry, exposure time, image quality, and sky brightness) using the suprême toolkit. [^4] https://github.com/LSSTDESC/supreme
The DC2 simulations have no throughput variation over the focal plane or between visits, so joint photometric and astrometric calibration across visits is not performed. Consequently, the standard passbands for the LSST filters are simply the total throughputs used in the simulations and were derived by the Rubin systems engineering team. [^5] https://github.com/lsst/throughputs/tree/DC2production We used these same throughputs to generate the stellar populations of simulated dwarf galaxies as described in Section 2.4. Furthermore, DC2 does not include bright stars ($r\lesssim 10$ mag), regions of high stellar density, or large interstellar extinction. To mitigate the impact of these limitations, we mask regions of the LSST footprint where the survey performance is expected to differ appreciably from DC2 when making predictions for the observable population of Milky Way satellite galaxies (Section 5).
Figure 1: The simulated LSST DESC DC2 footprint covers $\sim$300 deg2 of the high Galactic latitude sky. Maps of the ${\rm S/N}=5$ magnitude limit for point-like sources in the $g$-band (left) and $r$-band (right) show the uniformity and depth of the WFD portion of DC2. Variations in the depth come from the the observing strategy and simulated observing conditions. The median ${\rm S/N}=5$ magnitude limits for point-like sources are $g=27.0$ mag and $r=26.8$ mag.
@@ 2.2 Data Access
We accessed and analyzed the DC2 data through the RSP hosted on the Interim Data Facility (IDF) in the context of the first data preview (DP0.1) released in the period leading up to the start of LSST (Guy et al. 2023). The goals of DP0.1 were to serve as an early integration test of the LSST Science Pipelines with the RSP, and to enable members of the scientific community to begin early preparations for science with LSST. DP0.1 was hosted on the IDF and access was provided through the RSP, which includes a set of integrated web-based applications, services, and tools to query, visualize, subset, and analyze LSST data (O’Mullane et al. 2021). The $\sim$300 deg2 WFD data set from DC2 was ingested and adopted as the primary data set for DP0.1.666https://dp0-1.lsst.io/data-products-dp0-1 In addition to the simulated output catalogs, DP0.1 contains a truth matched catalog to access properties of the astronomical objects used as input to DC2. All analyses presented in this paper were performed using the RSP on the IDF, and a discussion of how we optimized our analysis to run on the RSP is provided in Appendix A.
@@ 2.3 Catalog Characterization
To quantify star/galaxy separation efficiency, we selected a $\sim$3 deg2 region of the DC2 WFD footprint centered on $({\rm R.A.,Dec.})=(62\,\deg,-37\,\deg)$. Within this region, we selected a set of well-measured objects with truth matches and $g$- and $r$-band magnitude error, $\sigma_{g,r}<0.2$ (S/N $\gtrsim 5$). We performed a simple star/galaxy separation based on the measured EXTENDEDNESS parameter (Bosch et al. 2018; LSST DESC et al. 2022). The EXTENDEDNESS is a boolean classification variable defined from the difference between the point spread function (PSF) magnitude, and the CModel magnitude for objects. Point-like sources are have $\texttt{EXTENDEDNESS}=0$, while extended galaxies have $\texttt{EXTENDEDNESS}=1$ (LSST DESC et al. 2022). To characterize the detection efficiency, we also tracked all stars in the region that were truth matched regardless of whether they passed our $\sigma_{g,r}<0.2$ threshold. We binned these populations in true $r$-band magnitude to derive the stellar detection efficiency, stellar classification efficiency, and the simultaneous combination of the two (Figure 2). We find a steep drop in stellar classification efficiency by $r\sim 26$ mag.
@@
To derive a model for the photometric uncertainty as a function of object magnitude and survey depth, we used the measured magnitude uncertainties of stars in combination with ${\rm S/N}=5$ magnitude limit depth maps generated from the input images (Figure 1). The magnitude limit maps were generated from the simulated observing properties (i.e., exposure time, exposure geometry, PSF FWHM, sky brightness) of the DC2 simulations using the suprême package. [^7] https://github.com/LSSTDESC/supreme/ We compute the median of the logarithm of the $r$-band magnitude uncertainty as a function of the difference between the true input $r$-band magnitude of the simulated object and the value of the magnitude limit map at the location of the object, $\Delta r=r-{\rm maglim}(r)$ (Figure 2). This model for the expected magnitude uncertainty as a function of magnitude and survey magnitude limit is applied to the simulated satellite member stars to generate realistic photometric scatter in magnitude and color (Section 2.4).
Figure 3: Simulated LSST color–magnitude diagram of stars associated with a simulated Milky Way satellite (red points; $M_{V}=-0.3$ mag, $D=91$ kpc) and the density of simulated objects from DESC DC2 that are classified as stars (gray background). The left panel uses a measured star/galaxy classification based on the EXTENDEDNESS parameter while the right panel assumes a perfect star/galaxy separation. The red solid line shows our isochrone selection region, and the dashed red line indicates the magnitude cut at $g<26$ mag and $r<26$ mag that is applied in our analysis. The population of faint ($g\gtrsim 25$ mag), blue ($g-r\lesssim 0.25$) background galaxies that are mis-classified as stars presents a significant contamination in searches for faint Milky Way satellites. We note that the sharp break in stellar density at $g-r\sim 0.25$ comes from the main sequence turn-off of simulated Milky Way halo stars in DC2.
@@ 2.4 Simulated Satellites
The simulated LSST bandpasses used for DC2 were integrated into the Padova CMD interface to generate photometry consistent with DC2.888http://stev.oapd.inaf.it/cgi-bin/cmd The initial masses of satellite member stars were drawn from a Chabrier (2001) initial mass function (IMF), which has been found to be a reasonable description of known satellite galaxies (Simon 2019). These initial masses were used to assign current absolute magnitudes from a Marigo et al. (2017) isochrone. When sampling from the IMF, the lower mass bound was set to the hydrogen-burning limit of $0.08M_{\odot}$ and the upper bound was set by the star with the largest initial mass in the evolved isochrone (post-asymptotic giant branch and white dwarf stars are excluded). Using the Marigo et al. (2017) isochrones, we transformed from initial stellar mass to current absolute magnitude in the $g$ and $r$ bands of the DC2 photometric system, and then to apparent magnitudes using the distance modulus of the simulated satellite. Detectability, stellar classification probability, photometric scatter, and photometric measurement uncertainties were assigned to each simulated star based on its true apparent magnitude and the DC2 survey limiting magnitude at the location that the star was injected following the procedures described in Section 2.3. In this way, our simulated stellar catalog mimics the DC2 catalog in detection efficiency, object (mis-)classification, and photometric uncertainty, and thus it can be combined with the DC2 catalog directly.
@@ 3 Analysis
To explore the impact of star/galaxy classification efficiency, we ran the analysis in two different configurations. In the first, the star/galaxy classification was performed using the measured EXTENDEDNESS parameter: objects with $\texttt{EXTENDEDNESS}=0$ were classified as stars, and those with $\texttt{EXTENDEDNESS}=1$ were classified as background galaxies. In the second configuration, the star/galaxy classification instead used the true object class from the simulation inputs. The difference in star/galaxy classification efficiency was also applied when generating simulated satellites—i.e., we generated two separate sets of satellite simulations applying the measured and perfect star/galaxy classification efficiency. Example color–magnitude diagrams of the objects passing these selections for one simulated satellite are shown in Figure 3. Downstream analysis procedures were the same for these two different star/galaxy separation procedures. We note that the sharp break in the stellar density at $g-r\sim 0.25$ (Figure 3) comes from the main sequence turn-off of Milky Way halo stars in DC2 generated by galfast (Jurić et al. 2008). Similar sharp features are seen in deep photometric data (e.g., Pieres et al. 2020) and other simulations of the LSST stellar population (e.g., Dal Tio et al. 2022).
@@ 6 Conclusions
While DC2 is the highest-fidelity simulation of LSST to date, there are several limitations in using it to derive the LSST survey selection function for Milky Way satellites. First, DC2 only covers a small region of the high-Galactic-latitude sky and does not capture the full range of inhomogeneities expected in the LSST WFD footprint (e.g., stellar density, interstellar extinction, and survey depth). To mitigate the impact on our predictions for the observable Milky Way satellite galaxy population, we explicitly mask regions of the sky where the LSST WFD performance is expected to differ significantly from DC2. Secondly, DC2 was simulated using an out-of-date model for the 5-year LSST survey cadence and environmental conditions. The median $g$- and $r$-band depths from DC2 are compatible with more recent simulations of the baseline LSST WFD strategy, and we restrict our predictions to the region of the LSST WFD footprint that matches the DC2 depth. Finally, our results come from injecting simulated stellar populations of Milky Way satellites into the DC2 catalogs. Recent studies have shown that catalog-level simulations give slightly higher sensitivity than full image-level simulation (e.g., Zhang et al. 2025); however, these differences are relatively small in the ranges of satellite size, luminosity, and distance that we study here.
[…]
There also exist novel ideas for improving star/galaxy separation efficiency using multi-wavelength observations and advanced image processing techniques (e.g., Kovács & Szapudi 2015; Sevilla-Noarbe et al. 2018; Muyskens et al. 2022; Bechtol et al. 2025). Satellite detection efficiency can also be improved by utilizing likelihood-based search algorithms (e.g., Bechtol et al. 2015; Drlica-Wagner et al. 2020). These algorithms construct a likelihood function from the product of Poisson probabilities to detect individual stars based upon their spatial positions, measured fluxes, photometric uncertainties, and the local imaging depth, given a model that includes a putative dwarf galaxy and empirical estimation of the local stellar field population. The combination of the simple matched-filter algorithm applied here with a likelihood-based search approach has been found to greatly reduce false positives (Drlica-Wagner et al. 2020).
@@
Software: Astropy (Astropy Collaboration et al. 2013, 2018), HealPy (Zonca et al. 2019) healsparse (Rykoff & Sánchez 2019), Matplotlib (Hunter 2007), Numpy (Harris et al. 2020), Pandas (The pandas development team 2020), Phalanx [^15] https://phalanx.lsst.io/, SciPy (Virtanen et al. 2020), simple (Bechtol et al. 2015), ugali (Bechtol et al. 2015; Drlica-Wagner et al. 2020)